# Open Your Ears to Take a Look: A State-of-the-Art Report on the Integration of Sonification and Visualization

### v1

@@ 3.10 Users and Goals
Just as with visualization-only designs, audiovisual display idioms are often developed toward specific users and end goals. The most commonly occurring goal for the papers in the database was data analysis. This could either be to present a dataset to the user or letting the user explore one or several datasets with the idiom, as described in subsection 3.2. The benefit of using visualization and sonification in this regard is that the user can gain different perspectives of the data through the two sensory modalities. For example, Alonso-Arevalo et al. [$*{}^{*}$12] created a multimodal interface for curve shapes and curvature, where the user can evaluate the quality of a three-dimensional shape by using both the visual and auditory, and in this case even their haptic perception. The curve shape was mapped to the fundamental frequency of different carrier sounds to offer different sound designs to convey the information.
@@
The most recurring pair of users and goals in the database was data analysis for domain experts, which most commonly would involve an audiovisual display idiom to convey data in a specific application domain in the natural sciences. Temor et al. [$*{}^{*}$21] and MacDonald et al. [$*{}^{*}$18] created an auditory complement to a scientific visualization of computational fluid dynamics, where the sonification aided in understanding the temporal changes in the visual animation. Gionfrida at al. [$*{}^{*}$16] suggest to combine the visualization of brainscans (in the context of Alzheimer’s dementia research) with a parameter mapping sonification that they call Triple-Tone Sonification. The design makes use of the fact that the human ear will perceive the sound of two or more very similar frequencies as "frequency beating," which is also employed in [MAFP19a]. For the purpose of process execution data analysis, Hildebrandt et al. [HAR16] demonstrated how incorporating sound could enhance the process of identifying anomalies or conducting root cause analysis for irregularities and errors. A design for a more general user case was created by North et al. [NSC16], which created a an idiom to convey Git version control data. The temporal nature of the data lent itself well to be sonified by sequentially going through the data, where an earcon is played when a commit occurs, and drum sounds are played whenever a conflict occurs in the data.
@@
The second most commonly occurring pair of users and goals was public engagement for the general public, where the use of an audiovisual display idiom has the potential to enable higher engagement by conveying information through both the visual and auditory modality. This would commonly be targeted towards popular science topics such as climate change [HK22, Bal15] and astronomy [HTHB22, TB23]. As previously mentioned in subsection 3.2, the work by Russo et al. [RS22] is an example of a sonification targeted towards the general public for public engagement, where the response of the resulting videos on social media were in part used as a metric for the public engagement of the design.
@@
Discussion: Regarding the successful design of audiovisual idioms, actually supporting their users in analyzing data, it is most relevant to include them in the design process. Both the visualization community and the sonification community have individually studied the relevance of including domain experts in their design process [SMM12, Gou17], and the same will be necessary for the integration of sonification and visualization. The co-author network shown in Figure 9 displays quite many domain experts being collaborators in the surveyed literature, which can be regarded as a promising sign in general.
@@ 4.1 Correlation Matrix Analysis
To understand potential relationships between different tags within our classification, we computed a correlation matrix, shown in Figure 8. We want to highlight some of the found correlations in the data. Note that overall most of the correlations are close to zero and are most likely not significant. Also, some of the higher correlation values that can be found in the matrix are very likely not significant, due to the low amount of data for these tags (e.g., the AP evaluation, which has been used in only two papers). Therefore, we only discuss correlations that are plausible and most relevant to the bigger picture.
Close to the main diagonal of the correlation matrix we highlighted four smaller matrices (1), showing an interesting phenomenon: Whenever the tagging of the papers was done individually for the visualization and the sonification parts, the two were mostly tagged with the same label. Hence, the visualizations and the sonification, in many cases, use the same reading levels [Ber83], the same search levels [Mun15], the same dataset types, and the same levels of measurement. This finding can motivate future research regarding the reading levels and the search levels. Studies could investigate the potentials of other distributions of the reading level onto the senses.
Marked with (2), we highlighted four phenomena concerning the purpose of an audiovisual idiom. Idioms designed to present rather than explore data are likely to also use the reading level “whole” for their sonification ($r=0.41$). On the other hand, idioms that are designed for exploration, likely use the group reading level ($r=0.3$). We see a correlation between the purpose of an idiom and its interactivity, with idioms for exploration more likely to be interactive ($r=0.23$), and idioms for presentation more likely not to be interactive ($r=-0.48$). Also, domain experts use more idioms for exploration ($r=0.44$), while the general public is more likely to be ‘just’ presented with data ($r=0.50$), going hand in hand with their general goals of data analysis ($r=0.52$) or public engagement ($r=0.50$).
Two more relationships with the interaction tag emerge form the correlation matrix (3). A sonification that uses the reading levels of “group” or “single,” is likely to offer user interaction ($r_{group}=0.45$ and $r_{single}=0.24$), while a sonification with the reading level of “whole” usually does not offer user interaction ($r_{whole}=-0.28$). Furthermore, designs that map data to the senses in a complementary way are more likely to offer interaction to their users ($r=0.32$).
@@ 4.2 Co-Author Network and Development Over Time
Figure 9: The co-author network, displaying 165 authors of the 57 papers in the database. The links represent papers written as a team of co-authors, and the colors display their primary disciplines. An interactive version of the network can be explored using the Orange workflow that is part of the supplemental material.
@@
As part of our meta analyses, we studied the co-author network within our database. Each node in the co-author network represents an author with a total of 165 nodes (in 48 teams), and a link connects two nodes if the corresponding authors have collaboratively contributed to a publication. The thickness of the links is proportional to the number of publications shared between the respective authors. We assign a color to each node using the primary discipline of the author. We classify each author, to the best of our abilities, into one of four disciplines: sonification, visualization, domain, and unclassified. For each author, we consider their main publication focus particularly around the time of their contribution to the works in our database, along with their background, education, and their role and input to the relevant papers. Authors classified under “domain” are recognized as domain experts and the category “unclassified” encompasses scenarios where the author’s primary discipline does not fall into any of the other three categories, the author is in the early stages of their career without a defined research focus, or that the information about the author is not available.
@@
The network offers several insights into the structure of the field. The majority of co-author teams are individual teams without interconnections. Only in four cases have authors collaborated with different (groups of) co-authors, usually because one person worked with different collaborators. This phenomenon differs from networks in other fields, such as the one described in a recent STAR on the use of embeddings in Visual Analytics [HWKK23], showing a network with clusters up to the size of 75 authors (note that the STAR on embeddings has 122 entries). We interpret the dominance of many disjoint groups as a sign for the rather young field and hope, with this STAR, to contribute to the future development of the field growing closer together. We identified 14 teams or individual authors contributing to the field that have only sonification experts on their team. Two of the teams consist exclusively of visualization researchers and 21 of the teams are collaborations between the communities, also including domain experts many times. The inclusion of domain experts and the quite large number of diverse groups also including domain experts can be considered a promising sign for the future development of the field.
Figure 10: The bar chart shows the number of papers published between the years 2011 to 2023. Generally, we see a rising trend in publications over the course of the last 12 years.
@@
During the years 2011 and 2023, overall, 57 papers were published. Figure 10 displays the majority of papers being published in the second half of this time span, pointing towards a generally increasing trend. The drop in publications in the year 2020 is likely a phenomenon related to the COVID-19 pandemic, being dominant worldwide especially during the year 2020.
@@ 5 Adjacent Topics
As mentioned earlier, three distinct topics are adjacent to the scope of our STAR: audiovisual idioms in the context of accessibility, real-time monitoring, and arts. Although the three areas are related, they might be better served by a different classification system than the one we used. Still, they are relevant and can be inspirational for our field. The following three sections present work from those adjacent fields that have been curated by our team of co-authors. The sections do not claim completeness or the usage of a systematic approach but are indented as a general introduction for the interested reader.
@@ 5.1 Accessibility
Even though visualization is one of the most common ways of communicating data, many visualizations are inaccessible to readers with visual impairments. Some studies have suggested the use of sonification to support readers with visual impairments. During the filtering and classification of the assembled literature, a category emerged exploring sonification for accessibility of visualization. The literature that was sorted into this category focused on the design and evaluation of sonification to improve accessibility of visualizations for visually impaired people. As the literature in this category was not exploring integrated audiovisual representations for data analysis in general, but rather using sonification as a support tool, these are not included in the systematic part of our STAR.
@@
Sonification for accessibility in this way often suggests sonification designs leaning towards auditory graphs. Auditory graphs are the auditory equivalent to visual representations, such as plots, graphs, and charts [HHN11], but instead of mapping data properties to line positions or the size of an area, they are mapped to various auditory parameters such as pitch or loudness. Auditory graphs are suggested to be useful as sensory substitution and as a means of presenting information when line of sight is not possible [SNH05], and using sound to represent data might make data analysis more accessible [WM10]. To give the reader a brief introduction to this category, here are a few recent studies that have explored sonification for accessibility in visualization.
@@
One study explored the accessibility of data visualizations presenting data to blind and visually impaired users [$*{}^{*}$23]. It was found that sonification might make discerning trends in the data more accessible, but also that the lack of experience in using sonification could lead to misinterpretations of the presented data. The study was conducted using screen readers, and the users often used alt-text information to support and validate what was heard through the sonification. Another study explored the use of natural sounds mapped to data visualizations, bar charts, and line charts [HEUHEB23]. The reason for using natural sounds was to support users without musical training in understanding the sonification and auditory graph, and it was found that these natural sounds could support the understanding of categorical data and were most useful for users without musical training. Infographics, the combination of visualization and text information, has been explored with an auditory-only approach [$*{}^{*}$22]. An interactive approach, infosonic, was explored to facilitate accessibility of data to blind and low-vision users, using spoken introduction and annotation, and non-speech sonification. The study shows that the sonification approach supported understanding and forming a mental image of the data. Even if only a few studies have been introduced in this text there are more studies on sonification and accessibility, and this is important for individuals who are blind or of low vision to have equitable access to data, news, and information.
@@
Often accessibility studies tend to blend sonification, i.e., non-speech sounds, with generated speech or screen readers. This makes the interpretation of the results from these studies somewhat challenging. The findings, i.e., the experiences and the understanding of the sonification, might be due to the sonification itself, or depending on other auditory stimuli used. Also, the visual representation, the user interaction, and user-interface are likewise of importance for the study outcome. In visual data analysis interaction is essential for exploring the data [CGM19], and similarly, for sonification to be a useful tool dynamic human interaction is necessary [HH04]. Nevertheless, sonification for accessibility of visualization for the visually impaired is an interesting approach. Some of these findings might suggest interesting approaches also for readers with normal or corrected to normal visual acuity in supporting visual perception with sonification [Rön19a, Rön19b], or reducing cognitive load [ZPR16] on the visual modality by sonification (see for example [MLS95]).
@@ 5.2 Monitoring
A comprehensive overview or a systematic literature review of audiovisual process monitoring is far beyond the scope of this report. Therefore, this section is intended to provide some insights into this research field and to distinguish the various scientific communities involved.
@@
The combination of visual and auditory displays for process monitoring purposes has been well-established in real-world applications and interfaces. The audiovisual representation of Morse code for an exchange of information serves as an early example of a multi-modal monitoring display and interface. In supervising and controlling the states of various devices and processes, audiovisual interfaces combine the advantages of both modes: volatile alerting sound signals enhance situation awareness, while visual cues provide additional information about the situation, enabling users to take action. Implementations of audiovisual interfaces are ubiquitous in both professional and day-life environments. For instance, consider the warning beep, which can be heard, when the temperature outside your car drops below 4°C, alerting you to the possibility of slippery road conditions. A quick glance at the visual display will provide you with the exact temperature and the opportunity to decide whether to take action or not.
@@
Besides the approaches to audiovisual data analysis described in the previous sections of this report, audiovisual monitoring interfaces play an integral role in various domains, including air traffic control [$*{}^{*}$23], control rooms [SFLD22, $*{}^{*}$12], anesthesia [$*{}^{*}$21, $*{}^{*}$22], neurology [$*{}^{*}$28], dermoscopy [$*{}^{*}$19], surgery [Zie23], network monitoring [Wor19a, $*{}^{*}$21], automotive [JDS15, $*{}^{*}$22], to name a few.
@@
During literature search for this STAR, several publications were labeled with the tag “monitoring”. After closer investigation, which involved excluding multiple publications written by identical authors on very similar topics, as well as publications addressing topics of accessibility and artistic media installations, only a relatively small representative subset of 14 findings qualifies for audiovisual process monitoring within the search criteria of this STAR.
@@
The reasons for this rather limited number of retrieved articles in this category are manifold. Firstly, most research on monitoring applications used for auditory feedback of human behavior in the fields of sports, therapy, or rehabilitation focuses on the design, impact, and evaluation of interfaces that utilize the non-intrusive attributes of auditory displays, enabling users to perceive information without any interfering actions, such as adjusting their head or body position to view a visual display [SJMT19, VRGM20, MNS16]. While some of these interfaces include visualizations at a low scale level, they hardly qualify as audiovisual monitoring devices in terms of providing a balanced contribution from both modalities.
@@
In industrial and surveillance contexts, which is another reason for the limited number of results in our search, factors such as cognitive workload, perception organization, situational awareness, alarm fatigue, and deafness play crucial roles in the successful implementation of infrastructures. However, these attributes are discussed in other scientific communities, including Human Factors and Applied Ergonomics, as well as several areas of medicine and health, particularly in anesthesia, rehabilitation, sports, and cognitive psychology. In addition to performance comparisons between auditory, visual, and audiovisual modalities [AG19], particular interest is given to the dual-task paradigm, which evaluates, for instance, the ability to identify an event requiring action in a secondary domain while simultaneously performing a primary task [$*{}^{*}$13, HHRM16, $*{}^{*}$15, NGJW14].
@@
(Non-)systematic literature reviews have been published in several of the mentioned areas [$*{}^{*}$17, KIK19, SMS21, HRM15]. To the best knowledge of the authors, no comprehensive STAR covering the entire spectrum of audiovisual or related multi-modal monitoring applications and approaches has been conducted thus far. However, there are some fundamental publications treating the specific attributes and criteria for research and development in the field, such as [Ibe20, Joh04, WS07].

### v2

@@ 3.10 Users and Goals
Just as with visualization-only designs, audiovisual display idioms are often developed toward specific users and end goals. The most commonly occurring goal for the papers in the database was “data analysis.” This could either be to present a dataset to the user or letting the user explore one or several datasets with the idiom, as described in subsection 3.2. The benefit of using visualization and sonification in this regard is that the user can gain different perspectives of the data through the two sensory modalities. For example, Alonso-Arevalo et al. [ASH12] created a multimodal interface for curve shapes and curvature, where the user can evaluate the quality of a three-dimensional shape by using both the visual and auditory, and in this case even their haptic perception. The curve shape was mapped to the fundamental frequency of different carrier sounds, and the idiom offered different sound designs to convey the information.
@@
The most recurring pair of users and goals in the database was data analysis for domain experts, which most commonly would involve an audiovisual display idiom to convey data in a specific application domain in the natural sciences. Temor et al. [TMN21] and MacDonald et al. [MNW18] created an auditory complement to a scientific visualization of computational fluid dynamics, where the sonification aided in understanding the temporal changes in the visual animation. Gionfrida et al. [GRK16] suggest combining the visualization of brain scans (in the context of Alzheimer’s dementia research) with a parameter mapping sonification that they call Triple-Tone Sonification. The design makes use of the fact that the human ear will perceive the sound of two or more very similar frequencies as "frequency beating," which is also employed in [MAFP19a]. For the purpose of process execution data analysis, Hildebrandt et al. [HAR16] demonstrated how incorporating sound could enhance the process of identifying anomalies or conducting root cause analysis for irregularities and errors. A design for a more general user case was created by North et al. [NSC16], which created a an idiom to convey Git version control data. The temporal nature of the data lent itself well to be sonified by sequentially going through the data, where an earcon is played when a commit occurs, and drum sounds are played whenever a conflict occurs in the data.
@@
The second most commonly occurring pair of users and goals was public engagement for the general public, where the use of an audiovisual display idiom has the potential to enable higher engagement by conveying information through both the visual and auditory modality. This would commonly be targeted towards popular science topics such as climate change [HK22, Bal15] and astronomy [HTHB22, TB23]. As previously mentioned in subsection 3.2, the work by Russo et al. [RS22] is an example of a sonification targeted towards the general public for public engagement, where the response of the resulting videos on social media was in part used as a metric for the public engagement of the design.
@@
Discussion: Regarding the successful design of audiovisual idioms that support their users in analyzing data, it is most relevant to include them in the design process. Both the visualization community and the sonification community have individually studied the relevance of including domain experts in their design process [SMM12, Gou17], and the same will be necessary for the integration of sonification and visualization. The co-author network shown in Figure 9 displays quite many domain experts being collaborators in the surveyed literature, which can be regarded as a promising sign in general.
@@ 4.1 Correlation Matrix Analysis
To understand potential relationships between different tags within our classification, we computed a correlation matrix, shown in Figure 8. We want to highlight some of the found correlations in the data out of which most are not surprising, nevertheless, they are not obvious and our data enables us to take such a meta-perspective.
Close to the main diagonal of the correlation matrix, we highlighted four smaller matrices (1), showing an interesting phenomenon: Whenever the tagging of the papers was done individually for the visualization and the sonification parts, the two were mostly tagged with the same label. Hence, the visualizations and the sonification, in many cases, use the same reading levels [Ber83], the same search levels [Mun15], the same dataset types, and the same levels of measurement. This finding can motivate future research regarding the reading levels and the search levels. Studies could investigate the potential of other distributions of the reading level onto the senses.
Marked with (2), we highlighted four phenomena concerning the purpose of an audiovisual idiom. Idioms designed to present rather than explore data are likely to also use the reading level “whole” for their sonification ($r=0.41$). On the other hand, idioms that are designed for exploration likely use the group reading level ($r=0.3$). We see a correlation between the purpose of an idiom and its interactivity, with idioms for exploration more likely to be interactive ($r=0.23$) and idioms for presentation more likely not to be interactive ($r=-0.48$). Also, domain experts use more idioms for exploration ($r=0.44$), while the general public is more likely to be ‘just’ presented with data ($r=0.50$), going hand in hand with their general goals of data analysis ($r=0.52$) or public engagement ($r=0.50$).
Two more relationships with the interaction tag emerge from the correlation matrix (3). A sonification that uses the reading levels of “group” or “single” is likely to offer user interaction ($r_{group}=0.45$ and $r_{single}=0.24$), while a sonification with the reading level of “whole” does not usually offer user interaction ($r_{whole}=-0.28$). Furthermore, designs that map data to the senses in a complementary way are more likely to offer interaction to their users ($r=0.32$).
Overall, most of the computed correlation values in the correlation matrix are close to zero. Also, some of the existing higher values are likely not significant due to the low amount of data for the respective tags (such as with the AP evaluation, which has been used in only two papers). Therefore, we only discussed three perspectives that have considerable correlation values, are plausible, and that we consider most relevant to the bigger picture.
@@ 4.2 Co-Author Network and Development Over Time
Figure 9: The co-author network, displaying 165 authors of the 57 papers in the database. The links represent papers written by a team of co-authors, and the colors display their primary disciplines. An interactive version of the network can be explored using the Orange workflow, which is part of the supplemental material.
@@
As part of our meta-analyses, we studied the co-author network within our database, providing a good picture of the state of the research community interested in audiovisual analytics. Each node in the co-author network represents an author with a total of 165 nodes (in 48 teams), and a link connects two nodes if the corresponding authors have collaboratively contributed to a publication. We assigned a color to each node using the primary discipline of the author and classified them, to the best of our abilities, into one of four disciplines: sonification, visualization, domain, and unclassified. For each author, we considered their main publication focus, particularly around the time of their contribution to the works in our database, along with their background, education, their role and input to the relevant papers. Authors classified under “domain” are recognized as domain experts, and the category “unclassified” encompasses scenarios where the author’s primary discipline does not fall into any of the other three categories, the author is in the early stages of their career without a defined research focus, or that the information about the author is not available.
@@
The network offers several insights into the structure of the field. The majority of co-author teams are individual teams without interconnections. Only in four cases have authors collaborated with different (groups of) co-authors, usually because one person worked with different collaborators. This phenomenon differs from networks in other fields, such as the one described in a recent STAR on the use of embeddings in visual analytics [HWKK23], showing a network with clusters up to the size of 75 authors (note that the STAR on embeddings has 122 entries). We interpret the dominance of many disjoint groups as a sign for the rather young field and hope, with this STAR, to contribute to the future development of the field growing closer together. We identified 14 teams or individual authors contributing to the corpus that consist of sonification experts exclusively. Two of the teams consist only of visualization researchers, and 21 of the teams are collaborations between the communities, including domain experts in many cases. The inclusion of domain experts and the quite large number of diverse groups also including domain experts can be considered a promising sign for the future development of the field.
@@
Our data also enables us to take a basic temporal perspective on the development of the field. During the years 2011 and 2023, overall, 57 papers were published. The majority of papers were published in the second half of this time span, pointing towards a generally increasing trend. A drop in publications in the year 2020 is likely a phenomenon related to the COVID-19 pandemic, being dominant worldwide, especially during the year 2020.
@@ 5 Adjacent Topics
As mentioned earlier, three distinct topics are adjacent to the scope of our STAR: audiovisual idioms in the context of accessibility, real-time monitoring, and arts. Although the three areas are related, they might be better served by a classification system different from the one we used. Still, they are relevant and can be inspirational for our field. The following three sections present work from those adjacent fields that have been curated by our team of co-authors. The sections do not claim completeness or the usage of a systematic approach but are intended as a general introduction for the interested reader.
@@ 5.1 Accessibility
Even though visualization is one of the most common ways of communicating data, many visualizations are inaccessible to readers with visual impairments. Some studies have suggested the use of sonification to support readers with visual impairments and a recent STAR on accessibility research in visualization identified 16 out of 56 papers that utilize sonification [KJRK21]. During the filtering and classification of our assembled literature, a category emerged exploring sonification for accessibility of visualization. The literature that was sorted into this category focused on the design and evaluation of sonification to improve the accessibility of visualizations for visually impaired people. As the literature in this category did not explore integrated audiovisual representations for data analysis in general but rather used sonification as a support technique, these are not included in the systematic part of our STAR.
@@
Sonification for accessibility in this way often suggests sonification designs leaning towards auditory graphs. Auditory graphs are the auditory equivalent to visual representations, such as plots, graphs, and charts [HHN11], but instead of mapping data properties to line positions or the size of an area, they are mapped to various auditory parameters such as pitch or loudness. Auditory graphs are suggested to be useful as sensory substitution and as a means of presenting information when line of sight is not possible [SNH05], and using sound to represent data makes data analysis more accessible [WM10]. To give the reader a brief introduction to this category, here are a few recent studies that have explored sonification for accessibility in visualization.
@@
One study explored the accessibility of data visualizations presenting data to blind and visually impaired users [FFSR23]. It was found that sonification might make discerning trends in the data more accessible, but also that the lack of experience in using sonification could lead to misinterpretations of the presented data. The study was conducted using screen readers, and the users often used alt-text information to support and validate what was heard through the sonification. Another study explored the use of natural sounds mapped to data visualizations, bar charts, and line charts [HEUHEB23]. The reason for using natural sounds was to support users without musical training in understanding the sonification and auditory graph, and it was found that these natural sounds could support the understanding of categorical data and were most useful for users without musical training. Infographics, the combination of visualization and text information, has been explored with an auditory-only approach [HGI22]. An interactive approach, infosonic, was explored to facilitate accessibility of data to blind and low-vision users, using spoken introduction and annotation, and non-speech sonification. The study shows that the sonification approach supported understanding and forming a mental image of the data.
@@
Often, accessibility studies tend to blend sonification, i.e., non-speech sounds, with generated speech or screen readers. This makes the interpretation of the results from these studies somewhat challenging. The findings, i.e., the experiences and the understanding of the sonification, might be due to the sonification itself or depending on other auditory stimuli used. Also, the visual representation, the user interaction, and the user interface are of importance to the study outcome. In visual data analysis, interaction is essential for exploring the data [CGM19], and similarly, for sonification to be a useful tool, dynamic human interaction is necessary [HH04]. Sonification for accessibility of visualization for visually impaired users is an exciting approach and some of these findings could suggest interesting approaches also for sighted users in supporting visual perception with sonification [Rön19a, Rön19b], or reducing cognitive load [ZPR16] on the visual modality by sonification (see for example [MLS95]).
@@ 5.2 Monitoring
A comprehensive overview or a systematic literature review of audiovisual process monitoring is far beyond the scope of this report. This section is intended to provide some insights into this research field and to distinguish the various scientific communities involved.
@@
The combination of visual and auditory displays for process monitoring purposes has been well-established in real-world applications and interfaces. The audiovisual representation of Morse code for an exchange of information serves as an early example of a multimodal monitoring display and interface. In supervising and controlling the states of various devices and processes, audiovisual interfaces combine the advantages of both modes: volatile alerting sound signals enhance situation awareness, while visual cues provide additional information about the situation, enabling users to take action. Implementations of audiovisual interfaces are ubiquitous in both professional and day-life environments. For instance, consider the warning beep, which can be heard when the temperature outside your car drops below 4°C, alerting you to the possibility of slippery road conditions. A quick glance at the visual display will provide you with the exact temperature and the opportunity to decide whether to take action or not.
@@
Besides the approaches to audiovisual data analysis described in the previous sections of this report, audiovisual monitoring interfaces play an integral role in various domains, including air traffic control [EBL23], control rooms [SFLD22, HHG12], anesthesia [AQH21, RMS22], neurology [LCS28], dermoscopy [WRK19], surgery [Zie23], network monitoring [Wor19a, AHvR21], automotive [JDS15, XWX22], to name a few.
@@
During the literature search for this STAR, several publications were labeled with the tag “monitoring.” After closer investigation, which involved excluding multiple publications written by identical authors on very similar topics, as well as publications addressing topics of accessibility and artistic media installations, only a relatively small representative subset of 14 findings qualifies for audiovisual process monitoring within the search criteria of this STAR.
@@
The reasons for this rather limited number of retrieved articles in this category are manifold. Firstly, most research on monitoring applications used for auditory feedback of human behavior in the fields of sports, therapy, or rehabilitation focuses on the design, impact, and evaluation of interfaces that utilize the non-intrusive attributes of auditory displays. Hence, they enable users to perceive information without any interfering actions, such as adjusting their head or body position to view a visual display [SJMT19, VRGM20, MNS16]. While some of these interfaces include visualizations at a low scale level, they hardly qualify as audiovisual monitoring devices in terms of providing a balanced contribution from both modalities.
@@
In industrial and surveillance contexts, factors such as cognitive workload, perception organization, situational awareness, alarm fatigue, and deafness play crucial roles in the successful implementation of infrastructures. However, these attributes are discussed in other scientific communities, including Human Factors and Applied Ergonomics, as well as several areas of medicine and health, particularly in anesthesia, rehabilitation, sports, and cognitive psychology. In addition to performance comparisons between auditory, visual, and audiovisual modalities [AG19], particular interest is given to the dual-task paradigm, which evaluates, for instance, the ability to identify an event requiring action in a secondary domain while simultaneously performing a primary task [LWP13, HHRM16, TML15, NGJW14].
@@
(Non-)systematic literature reviews have been published in several of the mentioned areas [WMY17, KIK19, SMS21, HRM15]. To the best knowledge of the authors, no comprehensive STAR covering the entire spectrum of audiovisual or related multimodal monitoring applications and approaches has been conducted thus far. However, there are some fundamental publications treating the specific attributes and criteria for research and development in the field, such as [Ibe20, Joh04, WS07].